Caleb sold a put option on Canadian dollars for $.05 per unit. The strike price was $.85,
and the spot rate at the time the option was exercised was $.92. Assume Caleb
immediately sold off the Canadian dollars received when the option was exercised. Also
assume that there are 50,000 units in a Canadian dollar option. What was Caleb’s net
profit on the put option?
1
Expert's answer
2021-08-10T06:43:31-0400

The holder of the call option, whoever Caleb has sold the option to will exercise it if "S-X\geq0", otherwise he/she will let the option expire.

In this case,

"S-X=\$0.92-\$0.85=\$0.07\ge0"so the holder of the option will exercise it. In that case,the option sellers' Caleb net profit per unit of Canadian dollar is calculated as follows:

Net profit per C"\$=" selling price of currency-Buying price of currency+premium on the option.

"=0.92-0.85+0.07=0.14"

Since each option contract contains 50000units of Canadian dollars, Net profit per option"=50000 units×0.14 =\$7000"